# Molecular Docking via Weighted Subgraph Isomorphism on Quantum Annealers

### v1

@@ 1 Introduction
Molecular Docking [1, 2] is an essential step in the drug discovery process that aims at calculating the optimal position and shape of a small molecule, i.e. the ligand, when it bound to a target protein [3]. The docking process can be divided into two main tasks. First, the shape complementarity search[4] that involves the detection of three-dimensional poses of the ligand inside the active site of the protein (usually called pocket), i.e. conformations, positions, and orientations. In the standard approach used for shape complementarity search, 3D representations of the molecules are manipulated according to their degree of freedom: rigid roto-translation and fragment rotations along the rotatable bonds [5]. Second the binding affinity affinity evaluation, i.e. the pose and ligand evaluation via a scoring function.
In recent years, the field of Quantum Computing [6] (QC) has undergone significant developments and quantum techniques have been explored lately for molecular docking approaches in different contexts. In particular, Banchi et al. [7] proposed the use of a technique called Gaussian Boson Sampling to address a rigid bodies formulation of molecular docking by mapping it to the problem of finding the maximum clique in a graph. Mensa et al [8] instead used a quantum-enhanced machine learning technique known as Quantum Kernel Estimation for Ligand Based Virtual Screening (LB-VS). Finally, a recent work [9] explored the possibility of supporting molecular docking by exploiting a quantum computing technique called Quantum Annealing [10, 11, 12] (QA) which can be used to solve optimization problems. The work focused mainly on the simple ligand expansion process, where the task was to find the torsional configuration that maximizes molecular volume, or equivalently, that maximizes the internal distances between atoms inside the molecule.
In the present work instead, the aim is to attack the entire shape complementarity search phase by formulating the problem including the protein pocket, and also in a way that can be more easily digested by an annealing approach. In particular, the standard methodology which is based on rigid roto-translation and fragment rotations is replaced with a completely new approach that is more suited to a Quadratic Unconstrained Binary Optimization (QUBO) formulation, i.e. the one natively used for optimization problems solved with QA.
@@
We aimed to develop a purely geometrical Molecular Docking approach and execute it on the latest quantum annealing hardware (D-Wave Advantage and 2000Q), in order to understand the capabilities of these devices and compare their performances with respect to classical methods like simulated annealing.
@@ 2.1 D-Wave Quantum Annealer
$\mathcal{H}_{QA}={A(s)}\sum_{i}\hat{\sigma}^{(i)}_{x}+{B(s)}H_{P}$ (3)
@@
$H_{P}=\sum_{i}h_{i}\hat{\sigma}^{(i)}_{z}+\sum_{i>j}J_{i,j}\hat{\sigma}^{(i)}_{z}\hat{\sigma}^{(j)}_{z}$ (4)
@@ 3.1 From Pocket to Space-Grid
Given a molecular pocket, docking points within the pocket are generated using methods derived from the literature [16, CAVIARPASS, 14, 15]. Docking points $v$ are used as vertices of a weighted graph $G_{grid}=\{v,e_{u,v},w_{u,v}\}$ where $e_{u,v}$ are edges connecting points $u$ and $v$ with associated weights $w_{u,v}$ defined as the Euclidean distance between docking points $u$ and $v$, i.e. $w_{u,v}=d(u,v)=|u-v|$. The constructed weighted graph $G_{grid}$ identifies a 3D space grid inside the pocket. It is possible to consider all the edges between pocket points, i.e. $G_{grid}$ is a complete graph, or only a restricted number of edges (e.g. connecting only those edges below a given threshold distance), in this case, $G_{grid}$ is not a complete graph. In the following, we will consider only the first case, where $G_{grid}$ is a complete graph. This is due to the fact that the complexity (i.e. number of linear and quadratic terms) of the QUBO formulation (that will be introduced later on) does not change varying the density of $G_{grid}$.
@@ 3.2 From Ligand to molecular graph
Given a Ligand, simplifications related to its chemical and geometrical properties are applied in the pre-processing phase:
@@
  • Fixed dihedral angles edges $e^{dih}_{i,j}$. Given the above construction involving $e^{bond}_{i,j}$ and $e^{angle}_{i,j}$, it may happen that $G_{mol}$ structure admits rotations around edges that are not actually rotatable. This means that it is necessary to add edges $e^{dih}_{i,j}$ with appropriate distance weights in order to fix the dihedral angle for edges that are not rotatable.
@@ 4 QUBO formulation
xi,i={1 if vertex iGmol gets mapped to vertex iGgrid
0otherwise
[…]
0otherwise
[…]
@@
$H_{1}=\sum_{i}(1-\sum_{i^{\prime}}x_{i,i^{\prime}})^{2}$ (6)
@@
$H_{2}=\sum_{i,j\in{G_{mol}}}\sum_{i^{\prime},j^{\prime}\notin{G_{grid}}}x_{i,i^{\prime}}x_{j,j^{\prime}}$ (7)
@@
$H_{iso}=H_{1}+H_{2}=\sum_{i}(1-\sum_{i^{\prime}}x_{i,i^{\prime}})^{2}+\sum_{i,j\in{G_{mol}}}\sum_{i^{\prime},j^{\prime}\notin{G_{grid}}}x_{i,i^{\prime}}x_{j,j^{\prime}}$ (8)
@@
$H_{opt}=\sum_{i,j\in{G_{mol}}}\sum_{i^{\prime},j^{\prime}\in{G_{grid}}}(w_{i,j}-w_{i^{\prime},j^{\prime}})^{2}x_{i,i^{\prime}}x_{j,j^{\prime}}$ (9)
@@
$\mathcal{H}_{qubo}=AH_{iso}+H_{opt}=A\sum_{i}(1-\sum_{i^{\prime}}x_{i,i^{\prime}})^{2}+A\sum_{i,j\in{G_{mol}}}\sum_{i^{\prime},j^{\prime}\notin{G_{grid}}}x_{i,i^{\prime}}x_{j,j^{\prime}}+B\sum_{i,j\in{G_{mol}}}\sum_{i^{\prime},j^{\prime}\in{G_{grid}}}(w_{i,j}-w_{i^{\prime},j^{\prime}})^{2}x_{i,i^{\prime}}x_{j,j^{\prime}}$ (10)
@@ 6.1 Ligand size $N_{mol}=5$, space-grid size $N_{grid}=10$
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ 6.2 Ligand size $N_{mol}=5$, space-grid size $N_{grid}=15$
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption

### v2

@@ 1 Introduction
In the early stages of drug discovery, computational methods play a critical role in reducing the use of expensive and time-consuming wet lab experiments. By using in-silico techniques, researchers can efficiently filter and prioritize potential drug candidates, significantly reducing the probability of failure in later stages of development. Among these computational approaches, molecular docking [1, 2] is an essential step in the drug discovery process as it aims to compute the optimal position and shape of a small molecule, or ligand, when it binds to a target protein [3]. The docking process can be divided into two main tasks. The first task is the shape complementarity search [4] that involves the detection of three-dimensional poses of the ligand inside the active site of the protein (usually called pocket). Meanwhile, the second task is the binding affinity evaluation, where the pose is evaluated through a scoring function.
Due to the computational intensity and resource demands of molecular docking, a variety of techniques have been developed to address these challenges. Furthermore, with the rapid advancements in Quantum Computing [5] (QC), more of these techniques are increasingly incorporating quantum computing to enhance the accuracy of docking results.
Quantum computing offers several paradigms for problem-solving, including Quantum Annealing, which leverages quantum tunneling and superposition for combinatorial optimization; Gate-Based Quantum Computing, which uses quantum gates to manipulate qubits; and a Hybrid Classical-Quantum Approach, where a parametric quantum circuit collaborates with a classical optimizer to fine-tune parameters. This paper focuses specifically on shape complementarity search in molecular docking using quantum annealers. In this context, the conventional approach of rigid roto-translation and fragment rotations [6] is replaced by framing the problem as a QUBO (Quadratic Unconstrained Binary Optimization) problem.
@@
We aimed to develop a purely geometrical Molecular Docking approach and execute it on quantum annealing hardware (D-Wave Advantage and 2000Q), to understand the capabilities of these devices and compare their performances with respect to classical methods like simulated annealing.
2 Main Related Works
An increasing number of approaches are now leveraging quantum computers to enhance the drug discovery process [11, 12, 13, 14, 15, 16, 10, 17]. Among these, some of the most prominent focus on addressing the Molecular Docking problem [11, 12, 13, 14].
One notable method involves reformulating the Molecular Docking problem as a matching problem, where the goal is to map a ligand onto a grid near the target pocket, with the objective function approximating the binding free energy:
$\Delta G_{bind}\approx\sum_{i=1}^{n}w_{i,s_{i}}\quad s_{i}=1,\dots,N$ (1)
Here, $n$ represents the number of vertices in the ligand graph, while $N$ indicates the number of vertices in the protein pocket grid. The weight $w_{i,s_{i}}$ is calculated based on the method used to estimate the binding free energy when ligand vertex $i$ gets mapped on the pocket grid vertex $s_{i}$. Two examples of these methods are Grid Point Matching (GPM) and Feature Atom Matching (FAM) [11]. This approach is particularly suited for Quantum Annealers, as it can be naturally translated into a Quadratic Unconstrained Binary Optimization (QUBO) problem.
Alternative techniques reformulate the Molecular Docking problem as a maximum weighted clique problem [12, 13]. In these methods, the ligand and protein are represented as graphs, where the nodes correspond to pharmacophore points - key determinants of a molecule’s pharmacological and biological interactions - and the edges are weighted based on the Euclidean distance between node pairs. These individual graphs are then combined into an interaction graph, where we want to identify the maximum weighted clique. The weights are determined by the pharmacophore point types associated with each vertex. Various techniques can be applied to solve this problem, including Gaussian Boson Sampling [12] and Variational Quantum Algorithms (VQAs) [13].
The Molecular Docking problem can be also solved using a hybrid classical-quantum approach when it is reframed into a Subgraph Isomorphism problem [14]. In this case, the objective is to identify the presence of a ligand graph within a protein graph. Both graphs are represented by their adjacency matrices, where $A$ corresponds to the protein graph with dimensions $N_{A}\text{x}N_{A}$, and $B$ corresponds to the ligand graph with dimensions $N_{B}\text{x}N_{B}$. The challenge involves finding a permutation matrix $P$ that minimizes the following cost function:
$\mathcal{L}_{C}(P;A,B)=||SPAP^{T}S^{T}-B||^{2}_{F}$ (2)
where $S=(\mathcal{I}_{N_{B}}|\mathcal{O}_{N_{B},N_{A}-N_{B}})$ and its action on a matrix $A$ (i.e. $SAS^{T}$) effectively selects the upper $N_{B}\text{x}N_{B}$ block of $A$. After embedding the problem into a quantum circuit, they utilize an Ansatz to represent the permutation matrices, parameterized by continuous rotation angles $\theta$. The algorithm alternates between iterations of the stochastic gradient descent (SGD) algorithm, which optimizes the parameters $\theta$ of the Ansatz, and a classical sampler that evaluates several solutions sampled from the superposition of states represented by the current parameters.
@@ 3.1 D-Wave Quantum Annealer
$\begin{split}\mathcal{H}_{QA}={A(s)}\sum_{i}\hat{\sigma}^{(i)}_{x}+{B(s)}H_{P}\end{split}$ (5)
@@
$\begin{split}H_{P}=\sum_{i}h_{i}\hat{\sigma}^{(i)}_{z}+\sum_{i>j}J_{i,j}\hat{\sigma}^{(i)}_{z}\hat{\sigma}^{(j)}_{z}\end{split}$ (6)
@@ 4 Molecular Docking Problem Definition
Molecular docking is a fundamental technique in structure-based drug discovery aimed at accurately positioning a ligand within the binding pocket of a protein. The binding pocket is defined as a cavity (empty space) within the protein structure that provides the spatial constraints necessary for ligand positioning. In the proposed approach, (i) the ligand is represented as a graph capturing either its full molecular structure or a simplified version, (ii) the protein pocket is modelled as a grid, and (iii) the docking process is formulated as a weighted subgraph isomorphism problem mapping the ligand graph onto a subset of the pocket grid.
@@ 4.1 From Pocket to Space-Grid
Given a protein pocket, docking points within the pocket are generated using methods derived from the literature [19, 7, 8, 9]. Docking points $v$ are used as vertices of a weighted graph $G_{grid}=\{v,e_{u,v},w_{u,v}\}$ where $e_{u,v}$ are edges connecting points $u$ and $v$ with associated weights $w_{u,v}$ defined as the Euclidean distance between docking points $u$ and $v$, i.e. $w_{u,v}=d(u,v)=|u-v|$. The constructed weighted graph $G_{grid}$ identifies a 3D space grid inside the pocket. It is possible to consider all the edges between pocket points, i.e. $G_{grid}$ is a complete graph, or only a restricted number of edges (e.g. connecting only those edges below a given threshold distance), in this case, $G_{grid}$ is not a complete graph. In the following, we will consider only the first case, where $G_{grid}$ is a complete graph. This is due to the fact that the complexity (i.e. number of linear and quadratic terms) of the QUBO formulation (that will be introduced later on) does not change varying the density of $G_{grid}$.
@@ 4.2 From Ligand to molecular graph
In the pre-processing phase, several simplifications are applied to ligands’ chemical and geometric properties to align them with the constraints of the current Quantum Computer, specifically in terms of qubits and their interconnections:
@@
  • Fixed dihedral angles edges $e^{dih}_{i,j}$. Based on the construction involving $e^{bond}_{i,j}$ and $e^{angle}_{i,j}$, it is possible that the $G_{mol}$ structure allows rotations around edges that, for chemical reasons, should not be rotatable - because the bond is not of a single type - or because a bond was transformed into a rotatable one due to the ligand approximations introduced. Consequently, it is necessary to add the edges $e^{dih}_{i,j}$ with appropriate distance weights to fix the dihedral angles for edges that are not rotatable.
@@ 5 QUBO formulation
$x_{i,i^{\prime}}=\begin{cases}1&\mbox{ if vertex $i\in G_{mol}$ gets mapped tovertex $i^{\prime}\in G_{grid}$ }\\ 0&\mbox{otherwise }\end{cases}$
@@
$\begin{split}H_{1}=\sum_{i}(1-\sum_{i^{\prime}}x_{i,i^{\prime}})^{2}\end{split}$ (8)
@@
$\begin{split}H_{2}=\sum_{i,j\in{G_{mol}}}\sum_{i^{\prime},j^{\prime}\notin{G_{grid}}}x_{i,i^{\prime}}x_{j,j^{\prime}}\end{split}$ (9)
@@
$\begin{split}H_{iso}=H_{1}+H_{2}=\sum_{i}(1-\sum_{i^{\prime}}x_{i,i^{\prime}})^{2}+\sum_{i,j\in{G_{mol}}}\sum_{i^{\prime},j^{\prime}\notin{G_{grid}}}x_{i,i^{\prime}}x_{j,j^{\prime}}\end{split}$ (10)
@@
$\begin{split}H_{opt}=\sum_{i,j\in{G_{mol}}}\sum_{i^{\prime},j^{\prime}\in{G_{grid}}}(w_{i,j}-w_{i^{\prime},j^{\prime}})^{2}x_{i,i^{\prime}}x_{j,j^{\prime}}\end{split}$ (11)
@@
$\begin{split}\mathcal{H}_{qubo}=AH_{iso}+H_{opt}=&A\sum_{i}(1-\sum_{i^{\prime}}x_{i,i^{\prime}})^{2}+A\sum_{i,j\in{G_{mol}}}\sum_{i^{\prime},j^{\prime}\notin{G_{grid}}}x_{i,i^{\prime}}x_{j,j^{\prime}}\\ &+B\sum_{i,j\in{G_{mol}}}\sum_{i^{\prime},j^{\prime}\in{G_{grid}}}(w_{i,j}-w_{i^{\prime},j^{\prime}})^{2}x_{i,i^{\prime}}x_{j,j^{\prime}}\end{split}$ (12)
@@ 8 Conclusions
This work has been supported in terms of resources by CINECA-ISCRA initiative, and by EuroCC-Italy.